### v1

# KAN4TSF: Are KAN and KAN-based models Effective for Time Series Forecasting?

Abstract
Time series forecasting is a crucial task that predicts the future values of variables based on historical data. Time series forecasting techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods. Although existing methods have made significant progress, they still suffer from two challenges. The mathematical theory of mainstream deep learning-based methods does not establish a clear relation between network sizes and fitting capabilities, and these methods often lack interpretability. To this end, we introduce the Kolmogorov-Arnold Network (KAN) into time series forecasting research, which has better mathematical properties and interpretability. First, we propose the Reversible Mixture of KAN experts (RMoK) model, which is a KAN-based model for time series forecasting. RMoK uses a mixture-of-experts structure to assign variables to KAN experts. Then, we compare performance, integration, and speed between RMoK and various baselines on real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases. And we find the relationship between temporal feature weights and data periodicity through visualization, which roughly explains RMoK’s mechanism. Thus, we conclude that KAN and KAN-based models (RMoK) are effective in time series forecasting. Code is available at KAN4TSF: https://github.com/2448845600/KAN4TSF.
1 Introduction
Time series forecasting (TSF) is the task of using historical data to predict future states of variables. This research area includes a broad scope of applications, such as financial investment, weather forecasting, traffic estimation, and health management Bi et al. (2023); Gao et al. (2023); Savcisens et al. (2024); Han et al. (2024a). The machine learning community’s progress has long inspired time series forecasting technology: the popularity of early statistical learning methods gave rise to SVR and ARMIA, while the development of deep learning introduced MLP and Transformer into time series forecasting. At present, various time series forecasting methods cover almost all deep learning network architectures, such as RNN, CNN, Transformer, and MLP Nie et al. (2023); Wu et al. (2023); Han et al. (2024b). The forecasting models derived from different network architectures have their own advantages in forecasting performance, running speed, and resource usage.
Although deep learning-based models have made notable progress in time series forecasting, there are still several challenges. The universal approximation theorem (UAT), which is the mathematical foundation of most mainstream forecasting models, cannot provide a guarantee on the necessary network sizes (depths and widths) to approximate a predetermined continuous function with specific accuracy. And this theory can only achieve an approximation rather than a representation. The limitations of UAT have become the sword of Damocles hanging over time series forecasting. Furthermore, the prediction mechanism of existing models is black-box, resulting in a lack of interpretability. These nontransparent methods are suspected of being suitable for tasks that require a low tolerance for errors, such as medicine, law and finance.
Kolmogorov-Arnold Network (KAN) Liu et al. (2024a), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a novel approach to solving the above challenges. On the one hand, KART proves that a multivariate continuous function can be represented as a combination of finite univariate continuous functions. This theorem establishes the relationship between network size and input shape under the premise of representation. On the other hand, KAN offers a pruning strategy that simplifies the trained KAN into a set of symbolic functions, enabling the analysis of specific modules’ mechanisms, thereby significantly enhancing the network’s interpretability. In addition, KAN’s function fitting idea is consistent with the properties of time series, such as periodicity and trend, which is conducive to embedding prior knowledge into the network structure and improving the performance of the network.
Despite being a relatively recent proposal, KAN, which employs a trainable 1D B-spline functions to convert incoming signals, has already sparked numerous efforts to improve or broaden its capabilities. Some studies propose KAN’s variants which replace the B-splines with Chebyshev polynomials SS (2024), wavelet functions Bozorgasl and Chen (2024), Jacobi polynomials Aghaei (2024), ReLU functions Qiu et al. (2024), etc., to accelerate training speed and improve network performance. Other studies introduce KAN with existing popular network structures for various applications. For example, ConvKAN Bodner et al. (2024) and GraphKAN Zhang and Zhang (2024); Xu et al. (2024) are proposed for image processing and graph processing. In summary, KANs have been extensively empirically studied in vision and language Azam and Akhtar (2024); Yu et al. (2024). However, existing studies lack a KAN-based model that considers time series domain knowledge, making it impossible to verify whether KAN is effective in time series forecasting.
To this end, we aim to propose a KAN-based model for the time series forecasting task and evaluate its effectiveness from four perspectives: performance, integration, speed, and interpretability. First, we propose the Reversible Mixture of KAN Experts model (RMoK), a KAN-based time series forecasting model that uses multiple KAN variants as experts and a gating network to adaptively assign variables to specific experts for prediction. RMoK is implemented as a single-layer network because we hope that it can have similar performance and better interpretability than existing methods. Then, we use a unified training and evaluation setting to compare the performance of RMoK and current popular baselines on seven real-world datasets. The experimental results show that RMoK achieves state-of-the-art (SOTA) performance in most cases. Subsequently, we conduct a comprehensive empirical study on KAN-based models, including the comparison between KAN and Linear, the effect of integrating KANs with the Transformer, and the speed of the KAN-based models. Finally, we discuss the interpretability of RMoK using the example of weather prediction. We visualize the weights of temporal features at different time steps in KAN and find the correlation between the weight distribution and the periodicity of the data.
To sum up, the contributions of this work include:
  • To the best of our knowledge, this is the first work that comprehensively discusses the effectiveness of the booming KANs for time series forecasting.
  • To validate our claims, we propose the Reversible Mixture of KAN Experts model, which uses a single layer of the mixture of KAN experts to keep a balance between performance and interpretability.
  • We fairly compare the performance between RMoK and baselines on seven real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases. And we also conduct a comprehensive empirical study on KAN-based models about integration and speed.
  • We mine the relationship between time feature weights and data periodicity through visualization, which roughly explains the mechanism of RMoK.
In summary, compared with the baselines in terms of performance, integration, speed, and interpretability, we conclude that KAN is effective in time series forecasting.
2 Problem Defintion
In multivariate time series forecasting, given historical data $\mathcal{X}=[X_{1},\cdots,X_{T}]\in\mathbb{R}^{T\times C}$, where $T$ is the time steps of historical data and $C$ is the number of variates. The time series forecasting task is to predict $\mathcal{Y}=[X_{T+1},\cdots,X_{T+P}]\in\mathbb{R}^{P\times C}$ during future $P$ time steps.
3 Related Work
3.1 Time-Series Forecasting Models
Figure 1: The computational process of Linear and KAN layers under a certain input and output dimension.
Although Transformer-based methods have almost become the standard in CV and NLP, various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently.
Transformer-based time series forecasting models have strong performance but high time and memory complexity. Informer Zhou et al. (2021) proposed ProbSparse self-attention to reduce the complexity from $\mathcal{O}(T^{2})$ to $\mathcal{O}(T\log T)$. Pyraformer Liu et al. (2022a) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$. PatchTST Nie et al. (2023) and Crossformer Zhang and Yan (2023) use the patch operation to reduce the number of input tokens, thereby reducing time complexity. The performance of early Multi-layer perception-based (MLP-based) models is generally weaker than transformer-based methods. However, NLinear Zeng et al. (2023) and RLinear Li et al. (2023) combine different normalization methods with a single-layer MLP, and achieve performance that exceeds Transformer-based models on some datasets with extremely low computational cost.
Recurrent neural networks (RNNs) are suitable for handling sequence data, making them a favorable choice for time series analysis. The size of the hidden state in RNN is independent of the input time series length, so recent RNN-based time series prediction models, such as SegRNN Lin et al. (2023) and WITRAN Jia et al. (2024), apply the RNN structure to time series forecasting with longer input. Convolutional neural networks (CNNs) are frequently used in time series forecasting models in the form of 1D convolution, such as ModernTCN donghao and wang xue (2024) and SCINet Liu et al. (2022b), but TimesNet Wu et al. (2023) takes a different approach by converting the 1D time series into a 2D matrix through the Fourier transform and then using 2D convolution for prediction. In addition to the above four types of network structures, there is also time series forecasting work based on new network architectures such as Mamba Dao and Gu (2024); Wang et al. (2024) and RWKV Peng et al. (2023); Hou and Yu (2024).
3.2 Kolmogorov-Arnold Network
The Kolmogorov-Arnold representation theorem (KART) is the mathematical foundation of the Kolmogorov-Arnold Network (KAN) Liu et al. (2024a), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem. We show the different between KAN and MLP in Figure 1. Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers KAN network can be represented as:
$\operatorname{KAN}(\mathbf{x})=\left(\boldsymbol{\Phi}_{L}\circ\boldsymbol{\Phi}_{L-1}\circ\cdots\circ\boldsymbol{\Phi}_{2}\circ\boldsymbol{\Phi}_{1}\right)\mathbf{x},$ (1)
where $\boldsymbol{\Phi}_{l},l\in[1,2,\cdots,L]$ is a KAN layer, and the output dimension of each KAN layer can be expressed as: $[n_{1},n_{2},\cdots,n_{L}]$. Therefore, the transform process of $j$-th feature in $l$-th layer can be formed as:
$\mathbf{x}_{l,j}=\sum_{i=1}^{n_{l-1}}\phi_{l-1,j,i}\left(\mathbf{x}_{l-1,i}\right),\quad j=1,\cdots,n_{l},$ (2)
where $\phi$ consists two parts: spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
$\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}\mathrm{Spline}(x),$ (3)
where $\mathrm{Spline}(\cdot)$ is a linear combination of B-spline functions $\mathrm{Spline}(x)=\sum_{i}c_{i}B_{i}(x)$.
Recent studies have proposed a large number of KAN variants to expand its application scope and improve its performance. Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials SS (2024) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields. Some work extends KAN to the fields of vision and graphics and proposes ConvKAN and GraphKAN Zhang and Zhang (2024); Bodner et al. (2024). Few works try to apply KAN to time series, but there is a lack of sufficient theoretical and experimental analysis Vaca-Rubio et al. (2024). This paper aims to analyze the effectiveness of KAN and its variants in time series forecasting, we use XKAN to describe these methods uniformly (for example, TaylorKAN is KAN with Taylor polynomials).
4 RMoK
4.1 Mixture of KAN Experts Layer
Figure 2: The structure of RMoK and MoK layer.
Time series in real-world scenarios frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time. Moreover, there are significant distribution discrepancies between variables in multivariate time series. This poses a significant challenge to time series forecasting techniques, inevitably impacting the KAN-based methods as well. Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing Linear and Transformer-based networks: KAN has many variants using different spline functions. Considering that the special spline function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
Following this idea, we propose the mixture of KAN experts (MoK) layer, which is a combination of KAN and mixture of experts (MoE). MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data. KAN and its variants only differ in the spline function in Equation 3, so we use $\mathcal{K}(\cdot)$ to represent these methods uniformly in this paper. Our proposed MoK layer with $N$ experts can be simple formed as:
$\mathbf{x}_{l+1}=\sum_{i=1}^{N}\mathcal{G}(\mathbf{x}_{l})_{i}\mathcal{K}_{i}(\mathbf{x}_{l}),$ (4)
where $\mathcal{G}(\cdot)$ is a gating network. This mixture of experts structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
The gating network is the key module of the MoK layer which is responsible for learning the weight of each expert from the input data. The softmax gating network $\mathcal{G}_{\mathrm{softmax}}$ which uses a softmax function and a learnable weight matrix $\mathbf{w}_{g}$ to schedule input data as:
$\mathcal{G}_{\mathrm{softmax}}(\mathbf{x})=\mathrm{Softmax}(\mathbf{x}\mathbf{w}_{g}).$ (5)
This gating network is popular due to its simple structure. However, it activates all experts once, resulting in low efficiency when there are a large number of experts. Therefore, we adopt the sparse gating network Shazeer et al. (2017) which only activates the best matching top-$k$ experts. It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retains experts with the highest $k$ values. The sparse gating network can be described as:
$\mathcal{G}_{\mathrm{sparse}}(\mathbf{x})=\mathrm{Softmax}(\mathrm{KeepTopK}(H(\mathbf{x}),k)),$ (6)
$H(\mathbf{x})=\mathbf{x}\mathbf{w}_{g}+\mathrm{Norm}(\mathrm{Softplus}(\mathbf{x}\mathbf{w}_{\mathrm{noise}})),$ (7)
where $\mathrm{Norm}(\cdot)$ is standard normalization.
Table 1: The statistics of the seven TSF datasets.
Dataset Variates Timesteps Granularity
ETTh1,h2 7 17,420 1 hour
ETTm1,m2 7 69,680 15 min
ECL 321 26,304 1 hour
Traffic 862 17,544 1 hour
Weather 21 52,696 10 min
Table 2: Overall performance. Both the best results of KAN-based models and baselines are highlighted in bold, and the best results of all methods are marked in red.
Dataset P KAN-based Transformer-based CNN-based Linear-based
RMoK-S RMoK-B PatchTST FEDFormer TimesNet SCINet RLinear DLinear
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.326 0.360 0.320 0.358 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.355 0.376 0.345 0.372
192 0.367 0.382 0.364 0.383 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.391 0.392 0.380 0.389
336 0.400 0.404 0.395 0.405 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.424 0.415 0.413 0.413
720 0.462 0.439 0.457 0.440 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.450 0.474 0.453
avg 0.389 0.396 0.384 0.397 0.387 0.400 0.448 0.452 0.400 0.406 0.485 0.481 0.414 0.407 0.403 0.407
ETTm2 96 0.176 0.261 0.176 0.261 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.182 0.265 0.193 0.292
192 0.244 0.305 0.240 0.302 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.246 0.304 0.284 0.362
336 0.306 0.346 0.299 0.342 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.307 0.342 0.369 0.427
720 0.405 0.404 0.397 0.401 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.407 0.398 0.554 0.522
avg 0.283 0.329 0.278 0.326 0.281 0.326 0.305 0.349 0.291 0.333 0.571 0.537 0.286 0.327 0.350 0.401
ETTh1 96 0.382 0.396 0.374 0.397 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.386 0.395 0.386 0.400
192 0.430 0.426 0.419 0.429 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.437 0.424 0.437 0.432
336 0.468 0.443 0.461 0.450 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.479 0.446 0.481 0.459
720 0.450 0.458 0.474 0.467 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.481 0.470 0.519 0.516
avg 0.433 0.431 0.432 0.436 0.469 0.454 0.440 0.460 0.458 0.450 0.747 0.647 0.446 0.434 0.456 0.452
ETTh2 96 0.313 0.357 0.301 0.353 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.288 0.338 0.333 0.387
192 0.397 0.409 0.379 0.405 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.374 0.390 0.477 0.476
336 0.441 0.446 0.432 0.446 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.415 0.426 0.594 0.541
720 0.453 0.461 0.446 0.463 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.420 0.440 0.831 0.657
avg 0.401 0.418 0.390 0.417 0.387 0.407 0.437 0.449 0.414 0.427 0.954 0.723 0.374 0.398 0.559 0.515
ECL 96 0.187 0.273 0.178 0.267 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.201 0.281 0.197 0.282
192 0.193 0.280 0.187 0.274 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.201 0.283 0.196 0.285
336 0.210 0.295 0.204 0.290 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.215 0.298 0.209 0.301
720 0.256 0.330 0.247 0.323 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.257 0.331 0.245 0.333
avg 0.211 0.294 0.204 0.288 0.205 0.290 0.214 0.327 0.192 0.295 0.268 0.365 0.219 0.298 0.212 0.300
Traffic 96 0.601 0.367 0.541 0.340 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.649 0.389 0.650 0.396
192 0.576 0.354 0.529 0.330 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.601 0.366 0.598 0.370
336 0.588 0.359 0.545 0.334 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.609 0.369 0.605 0.373
720 0.625 0.376 0.580 0.351 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.647 0.387 0.645 0.394
avg 0.597 0.364 0.549 0.339 0.481 0.304 0.610 0.376 0.620 0.336 0.804 0.509 0.626 0.378 0.625 0.383
Weather 96 0.175 0.225 0.171 0.221 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.192 0.232 0.196 0.255
192 0.224 0.267 0.220 0.263 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.240 0.271 0.237 0.296
336 0.284 0.308 0.277 0.302 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.292 0.307 0.283 0.335
720 0.366 0.358 0.360 0.354 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.364 0.353 0.345 0.381
avg 0.262 0.290 0.257 0.285 0.259 0.281 0.309 0.360 0.259 0.287 0.292 0.363 0.272 0.291 0.265 0.317
Top-1 Count 31 23 7 12
Table 3: The performance of KAN-based models vs. Linear-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
with MoE w/o MoE
RMoK-S RMoL-S RWavKAN RTaylorKAN RLinear
Dataset P MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
96 0.326 0.360 0.335 0.367 0.341 0.370 0.340 0.368 0.355 0.376
192 0.367 0.382 0.374 0.385 0.374 0.384 0.380 0.386 0.391 0.392
336 0.400 0.404 0.407 0.406 0.406 0.405 0.412 0.406 0.424 0.415
720 0.462 0.439 0.470 0.440 0.465 0.438 0.474 0.439 0.487 0.450
ETTm1 avg 0.389 0.396 0.397 0.400 0.397 0.399 0.402 0.400 0.414 0.407
96 0.382 0.396 0.382 0.392 0.396 0.402 0.388 0.398 0.386 0.395
192 0.430 0.426 0.436 0.423 0.439 0.430 0.438 0.424 0.437 0.424
336 0.468 0.443 0.487 0.449 0.479 0.449 0.477 0.442 0.479 0.446
720 0.450 0.458 0.486 0.472 0.464 0.461 0.477 0.461 0.481 0.470
ETTh1 avg 0.433 0.431 0.448 0.434 0.444 0.435 0.445 0.431 0.446 0.434
96 0.175 0.225 0.174 0.223 0.180 0.228 0.182 0.232 0.192 0.232
192 0.224 0.267 0.226 0.267 0.227 0.268 0.235 0.274 0.240 0.271
336 0.284 0.308 0.286 0.308 0.284 0.308 0.291 0.312 0.292 0.307
720 0.366 0.358 0.367 0.358 0.368 0.360 0.370 0.360 0.364 0.353
Weather avg 0.262 0.290 0.263 0.289 0.264 0.291 0.269 0.294 0.272 0.291
Table 4: The performance of KAN within iTransformer. The best results of all methods are highlighted in bold. KAN, JacobKAN, and TaylorKAN are KANs with B-splines functions, Jacobi polynomials, and Taylor polynomials.
Dataset P iTransformer +KAN +JacobKAN +TaylorKAN +MoK (Ours)
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.341 0.373 0.689 0.547 0.340 0.373 0.331 0.365 0.329 0.364
192 0.388 0.397 0.707 0.556 0.382 0.395 0.379 0.392 0.376 0.390
336 0.426 0.420 0.720 0.565 0.418 0.419 0.417 0.416 0.412 0.415
720 0.498 0.458 0.746 0.582 0.487 0.456 0.484 0.450 0.478 0.453
avg 0.413 0.412 0.716 0.562 0.407 0.411 0.403 0.406 0.399 0.406
ETTh1 96 0.390 0.404 0.696 0.557 0.401 0.414 0.397 0.409 0.393 0.409
192 0.445 0.436 0.716 0.572 0.454 0.445 0.443 0.436 0.442 0.441
336 0.489 0.457 0.721 0.583 0.499 0.469 0.491 0.461 0.476 0.458
720 0.505 0.487 0.731 0.607 0.521 0.499 0.510 0.488 0.494 0.485
avg 0.457 0.446 0.716 0.580 0.469 0.457 0.460 0.448 0.451 0.448
4.2 Reversible Mixture of KAN Experts Model
We can get a sophisticated KAN-based model by stacking multiple KANs or replacing the linear layers of the existing models with KANs. However, we try to design a simple KAN-based model that is easy to analyze while achieving comparable performance to the most state-of-the-art time series forecasting methods.
Inspired by several successful single-layer methods Li et al. (2023); Zeng et al. (2023), we propose a simple, effective and interpretable KAN-based model, Reversible Mixture of KAN Experts Network (RMoK), which uses RevIN Kim et al. (2022) and single MoK layer, as shown in Figure 2. First, $\text{RevIN}^{+}$ (the normalization operation of RevIN) uses a learnable affine transformation to normalize the input time series of each variable. Then, the MoK layer obtains the prediction results based on the normalized time series features. Finally, the prediction results are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ (the denormalization operation of RevIN).
During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts. Following the previous work Shazeer et al. (2017), we add an load balancing loss function to encourage experts have equal importance. First, we count the weight of experts as loads, and calcuate the square of the coefficient of variation of the load values as additional loss:
$\mathrm{L}_{\mathrm{load-balancing}}=\mathrm{CV}(\mathrm{loads})^{2}.$ (8)
And the total loss function is the sum of the prediction loss and load balancing loss with weight $w_{l}$:
$\mathrm{L}=\mathrm{MSE}(\mathcal{Y},\hat{\mathcal{Y}})+w_{l}\cdot\mathrm{L}_{\mathrm{load-balancing}}.$ (9)
5 Experiments
In this section, we experimentally validate the effectiveness of our proposed KAN-based models, RMoK, on various time series forecasting benchmarks. Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, the impact of integrating the KAN and MoK layers into other network structures (Transformer), and the interpretability of RMoK.
5.1 Experimental Settings
Dataset.
We conduct extensive experiments on seven widely-used real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al. (2021), ECL, Traffic and Weather Lai et al. (2018), whose statistical information is shown in Table 1. We follow the same data processing operations used in TimesNet Wu et al. (2023), where the training, validation, and testing sets are divided according to chronological order.
Evaluation Metrics
Following previous works Zhou et al. (2021); Wu et al. (2023), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
Baselines
We select 6 well-acknowledged forecasting models as our baselines, including (1) Transformer-based methods: PatchTST Nie et al. (2023) and FEDformer Zhou et al. (2022); (2) CNN-based methods: TimesNet Wu et al. (2023) and SCINet Liu et al. (2022b); (3) Linear-based methods: RLinear Li et al. (2023) and DLinear Zeng et al. (2023).
Table 5: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1. Training batch size is 64 and infer batch size is 1.
Model Param Train (it/s) Infer (it/s)
DLinear 139 K 102.51 360.93
RLinear 69.9 K 92.11 345.04
iTransformer 3.6 M 51.03 188.79
PatchTST 7.6 M 43.92 193.58
RKAN 1.2 M 5.94 193.25
RWKAN 277 K 70.39 308.69
RTKAN 208 K 77.87 316.94
RJKAN (d=4) 345 K 76.77 280.06
RJKAN (d=6) 483 K 58.96 254.65
RMoK-S 972 K 34.86 190.57
5.2 Can KAN-based Models get SOTA Performance?
In this section, we conduct extensive experiments to compare the forecasting performance of KAN-based models with advanced baselines.
We use two versions of our proposed RMoK: RMoK-S (the small version, which has four experts), and RMoK-B (the base version, which has eight experts). We compare RMoKs with six popular Transformer-based, CNN-based and Linear-based baselines on seven benchmarks. The experimental results are shown in Table 2, where both the best results of KAN-based models and baselines are highlighted in bold and the best results of all models are marked in red. The results of baselines come from published papers Wu et al. (2023), and the results of RMoKs are average of four times experiments with fixed seed in [0, 1, 2, 3]. Surprisingly, RMoK achieves the best results in most cases. Considering that RMoK is a simple single-layer method that does not model correlations among variates, this empirical finding adequately demonstrates that KAN-based models are effective in the TSF task.
Specifically, we can divide these seven datasets into two groups based on the number of variables. On the four ETT subsets with fewer variables, RMoK outperforms the baselines most, and RMoK-S and RMoK-B exhibit their own strengths under different forecasting length $P$. On the Weather, ECL, and Traffic datasets with more variables, RMoK-B is significantly better than RMoK-S, indicating that the mixture of experts approach is suitable for dealing with a large number of variables. In addition, due to the intricate spatiotemporal correlation among various variables in Traffic dataset, the Transformer-based method PatchTST achieves the best results, whereas our proposed simple RMoK achieves suboptimal results that far exceeded other baselines.
5.3 Does KAN Outperform Linear?
In this section, we conduct ablation experiments to compare KAN-based models with Linear-based models in three time series forecasting datasets.
For fair comparison, we replace the four KAN experts in RMoK-S with four Linear experts to obtain a Linear-based baseline with the mixture of expert structure, naming RMoL-S. And we replace the entire MoK layer with a single KAN or Linear to obtain RWavKAN, RTaylorKAN and RLinear to analyze the performance of KAN variants and Linear on the TSF task. The experimental results are reported in Table 3, where all the results are average of 4 times experiments and the best results are highlighted in bold. We can conclude three useful empirical experiences. First, the KAN-based model outperforms the linear-based model in most cases. We speculate this phenomenon is due to KAN’s function representation idea, which is more efficient to capture the periodicity and trend in time series. Second, the mixture of expert structures is applicable to both KAN and Linear, which should be attributed to the fact that the gating network assigns variables to specific experts. Third, the performance of KAN-based models is affected by the specific function, which may be related to the intrinsic distribution of the time series.
Figure 3: Heatmap of the expert loads in the RMoK model on the Weather dataset. The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all test samples.
5.4 Can KAN be Integrated into Other Methods?
In this section, we verify whether KAN can be integrated into the existing time series forecasting models as a plug-in to improve performance. We select iTransformer Liu et al. (2024b) as baseline and replace the linear projections of all the attention modules with different KAN variants and MoK layer. The experiments are conducted on the ETT datasets. For fair comparison, we set the same model hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy, and repeat four times to report the average results in Table 4. While iTransformer with various KANs performs well only in ETTm1, iTransformer with MoK achieves the best performance in most cases on both ETTh1 and ETTm1. These experimental findings demonstrate that MoK is a successful form for integrating KAN into Transformer-based methods.
(a) Data visualization
(b) The weight of each input feature.
Figure 4: Visualization of KAN learning the periodicity in the data.
5.5 Are KAN-based Models Efficiency?
We report the size of model parameters, the training and infering speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720. We implement all methods in a unify code library with PyTorch Paszke et al. (2019), and the testing platform is a GPU server with NVIDIA A100 80GB GPUs. We set training batch size to 64, infer batch size to 1. The KAN, WKAN, TKAN, JAKN represent KANs with B-splines functions, Wavelet function (transform is mexican hat), Taylor polynomials (order=3), and Jacob polynomials (d is degree which sets to 4 or 6). As shown in Table 5, the running speed of KANs are affected by the specific implementation. KAN variant with Taylor polynomials can achieve close running efficiency to the Linear. With future hardware optimization, the efficiency of KANs will still be improved.
5.6 Are KAN-based Models Interpretable?
In this section, we aim to analyze the interpretability of RMoK for the time series forecasting task. First, we generate a heatmap to visualize the outputs of the gating network, which decide variable-expert assignment. This visualization demonstrates how the RMoK simplifies the multivariate time series forecasting task into multiple univariate forecasting subtasks. Then, we analyze what knowledge RMoK learns from real-world time-varying systems in each subtask.
Our proposed RMoK model uses a single MoK layer, which consists of gating network and KAN experts. The gating network outputs the matching score between variables and experts based on the input feature, and the top-$k$ matching experts are selected for each variable to predict the future state. We train an RMoK-S model with 4 experts on Weather dataset with 21 variables, then count the top-1 scores of all samples in the test set and generate a heatmap in Figure 3, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all test samples. From this figure, although the matching score is related to the time-varying input data, the variable is still closely related to the specific expert. Therefore, we roughly approximate that the trained RMoK is a linear combination of several KANs, and RMoK treats the multivariate forecasting task as multiple univariate forecasting tasks. Although the above process is crude, this simplification helps us analyze the explainability of RMoK in complex real-world scenarios.
After simplification, we try to analyze why KAN is effective in time series forecasting from the perspective of univariate and single-expert subtask. We use the temperature variable from Weather dataset which collects data every 10 minutes to train the KAN with B-splines function as forecasting model. We decompose and visualize data according to the trend, season, and resid terms through statsmodels  [^1] https://www.statsmodels.org/stable/index.html library. As shown in Figure 4, the temperature series has obvious daily periodicity, which is consistent with life experience. Then, we input the complete daily period (the past 144 time steps) data to predict the state of the next time step. We visualize the weight of each feature dimension of the trained RMoK. As shown in the Figure 4(b), there are three peaks in the feature weight, which are near 0, 72 and 144 time steps. 0 represents the temperature at the same time step of the previous day’s period, 144 represents the temperature at the adjacent moment, and 70 represents half a period. These three weight peaks correspond to the three zero points of the cosine function in one period. Finally, we conclude that RMoK can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
6 Conclusion
This work discusses the effectiveness of KANs in time series forecasting. Due to various variants with different spline functions of KAN, we propose a single-layer mixture of KAN experts model (RMoK) to alleviate the distribution variation in time series. We experimentally compared KAN with existing baseline methods of various network architectures in terms of performance, integration, efficiency, and interpretability. Experimental results on seven real-world datasets show that RMoK performs in most metrics, it is sufficient to conclude that KAN and KAN-based models are effective in the time series forecasting task, and KAN can gain a place in the increasingly fierce model structure competition. The single-layer KAN-based model proposed in this work is only a preliminary attempt to introduce KAN into time series forecasting. We hope that our work can help future studies improve the performance and interpretability of KAN-based models.
Appendix A Implementation Details
A.1 Datasets
We conduct extensive experiments on seven real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al. [2021], ECL, Traffic and Weather Lai et al. [2018], whose detail information is shown in Table 6. We follow the same data processing operations used in TimesNet Wu et al. [2023], where the training, validation, and testing sets are divided according to chronological order. The datasets we used cover multiple domains (electricity, transportation, and weather), which are sufficient to verify the generalizability of our method.
Table 6: The information of the seven time series forecasting datasets.
Dataset Variates Timesteps Granularity Domain
ETTh1,h2 7 17,420 1 hour Electricity
ETTm1,m2 7 69,680 15 min Electricity
ECL 321 26,304 1 hour Electricity
Traffic 862 17,544 1 hour Transportation
Weather 21 52,696 10 min Weather
A.2 Baselines
To comprehensively compare KAN with other baselines under different network structures, we select six well-known forecasting models as our baselines, including (1) Transformer-based methods: PatchTST and FEDformer; (2) CNN-based methods: TimesNet and SCINet; (3) Linear-based methods: RLinear and DLinear. We report the details of these baselines in Table 7.
Table 7: The information of baselines.
Name Type Cite
PatchTST Transformer Nie et al. [2023]
FEDformer Transformer Zhou et al. [2022]
TimesNet CNN Wu et al. [2023]
SCINet CNN Liu et al. [2022b]
RLinear Linear Li et al. [2023]
DLinear Linear Zeng et al. [2023]
A.3 Experimental Settings
We implement our method and baselines in a unify code library with PyTorch  [^2] https://github.com/pytorch/pytorch and Pytorch-lightning  [^3] https://github.com/Lightning-AI/pytorch-lightning (codes will be released after acceptance), and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs. To prevent the sample drop in the test phase where the samples of the last batch are discarded, we set the test batch size to 1. Following previous works Zhou et al. [2021], Wu et al. [2023], we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
$\operatorname{MAE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left|x_{i}-\hat{x}_{i}\right|,$ (10)
$\operatorname{MSE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left(x_{i}-\hat{x}_{i}\right)^{2},$
where $x_{i}$ denotes the $i$-th ground truth, $\hat{x_{i}}$ represents the $i$-th predicted values, and $N$ represents the number of testing samples.
Appendix B Experimental Analysis
We report hyperparameter sensitivity and performance robustness experimental results to further evaluate the effectiveness of KAN-based modes in time series forecasting.
B.1 Hyperparameter Sensitivity
We conduct hyperparameter experiments to compare the performance between MoK and KAN layers. Results are shown in Table 8, where WavKAN, TaylorKAN and JacobKAN are KANs with wavelet functions, Taylor polynomials, and Jacobi polynomials. The type of spline function affects KAN’s ability to model time series, but there is currently a lack of theoretical guidance on how to choose the best suitable spline function, which can only be determined through experimental methods. Thus, MoK becomes a best practice for KAN, which does not require a lot of experiments to select the optimal spline function. This cocktail-like solution achieves better performance than a single KAN layer in most cases and achieves overall performance improvements on datasets with a large number of variables (such as Weather, Traffic, ECL).
Table 8: The performance of KAN-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
Setting with MoE w/o MoE
MoK-S MoK-B RWavKAN RTaylorKAN RJacobKAN
MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1 96 0.326 0.360 0.320 0.358 0.341 0.370 0.340 0.368 0.353 0.376
192 0.367 0.382 0.364 0.383 0.374 0.384 0.380 0.386 0.386 0.391
336 0.400 0.404 0.395 0.405 0.406 0.405 0.412 0.406 0.413 0.410
720 0.462 0.439 0.457 0.440 0.465 0.438 0.474 0.439 0.465 0.442
avg 0.389 0.396 0.384 0.397 0.397 0.399 0.402 0.400 0.404 0.405
ETTm2 96 0.176 0.261 0.176 0.261 0.179 0.265 0.181 0.264 0.177 0.263
192 0.244 0.305 0.240 0.302 0.247 0.309 0.247 0.306 0.246 0.308
336 0.306 0.346 0.299 0.342 0.311 0.349 0.308 0.345 0.309 0.348
720 0.405 0.404 0.397 0.401 0.412 0.408 0.405 0.401 0.406 0.403
avg 0.283 0.329 0.278 0.326 0.287 0.333 0.285 0.329 0.284 0.331
ETTh1 96 0.382 0.396 0.374 0.397 0.396 0.402 0.388 0.398 0.389 0.411
192 0.430 0.426 0.419 0.429 0.439 0.430 0.438 0.424 0.434 0.437
336 0.468 0.443 0.461 0.450 0.479 0.449 0.477 0.442 0.464 0.451
720 0.450 0.458 0.474 0.467 0.464 0.461 0.477 0.461 0.452 0.460
avg 0.433 0.431 0.432 0.436 0.444 0.435 0.445 0.431 0.435 0.440
ETTh2 96 0.313 0.357 0.301 0.353 0.321 0.358 0.297 0.344 0.310 0.352
192 0.397 0.409 0.379 0.405 0.407 0.413 0.387 0.401 0.396 0.405
336 0.441 0.446 0.432 0.446 0.452 0.450 0.434 0.441 0.437 0.441
720 0.453 0.461 0.446 0.463 0.487 0.477 0.448 0.458 0.445 0.456
avg 0.401 0.418 0.390 0.417 0.417 0.424 0.392 0.411 0.397 0.414
ECL 96 0.187 0.273 0.178 0.267 0.202 0.284 0.203 0.282 0.201 0.284
192 0.193 0.280 0.187 0.274 0.203 0.285 0.205 0.285 0.200 0.285
336 0.210 0.295 0.204 0.290 0.218 0.299 0.221 0.300 0.216 0.301
720 0.256 0.330 0.247 0.323 0.262 0.332 0.265 0.333 0.260 0.333
avg 0.211 0.294 0.204 0.288 0.221 0.300 0.224 0.300 0.220 0.301
Weather 96 0.175 0.225 0.171 0.221 0.180 0.228 0.182 0.232 0.178 0.230
192 0.224 0.267 0.220 0.263 0.227 0.268 0.235 0.274 0.229 0.271
336 0.284 0.308 0.277 0.302 0.284 0.308 0.291 0.312 0.288 0.310
720 0.366 0.358 0.360 0.354 0.368 0.360 0.370 0.360 0.368 0.359
avg 0.262 0.290 0.257 0.285 0.264 0.291 0.269 0.294 0.266 0.292

### v2

@@
-KAN4TSF: Are KAN and KAN-based models Effective for Time Series Forecasting?
+Are KANs Effective for Multivariate Time Series Forecasting?
 Abstract
-Time series forecasting is a crucial task that predicts the future values of variables based on historical data.
-Time series forecasting techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods.
-Although existing methods have made significant progress, they still suffer from two challenges.
-The mathematical theory of mainstream deep learning-based methods does not establish a clear relation between network sizes and fitting capabilities, and these methods often lack interpretability.
-To this end, we introduce the Kolmogorov-Arnold Network (KAN) into time series forecasting research, which has better mathematical properties and interpretability.
-First, we propose the Reversible Mixture of KAN experts (RMoK) model, which is a KAN-based model for time series forecasting.
-RMoK uses a mixture-of-experts structure to assign variables to KAN experts.
-Then, we compare performance, integration, and speed between RMoK and various baselines on real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases.
-And we find the relationship between temporal feature weights and data periodicity through visualization, which roughly explains RMoK’s mechanism.
-Thus, we conclude that KAN and KAN-based models (RMoK) are effective in time series forecasting.
-Code is available at KAN4TSF: https://github.com/2448845600/KAN4TSF.
+Multivariate time series forecasting is a crucial task that predicts the future states based on historical inputs.
+Related techniques have been developing in parallel with the machine learning community, from early statistical learning methods to current deep learning methods.
+Despite their significant advancements, existing methods continue to struggle with the challenge of inadequate interpretability.
+The rise of the Kolmogorov-Arnold Network (KAN) provides a new perspective to solve this challenge, but current work has not yet concluded whether KAN is effective in time series forecasting tasks.
+In this paper, we aim to evaluate the effectiveness of KANs in time-series forecasting from the perspectives of performance, integrability, efficiency, and interpretability.
+To this end, we propose the Multi-layer Mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions.
+The core module of MMK is the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts.
+Then, we explore some useful experimental strategies to deal with the issues in the training stage.
+Finally, we compare MMK and various baselines on seven datasets.
+Extensive experimental and visualization results demonstrate that KANs are effective in multivariate time series forecasting.
+Code is available at: https://github.com/2448845600/EasyTSF.
 1 Introduction
-Time series forecasting (TSF) is the task of using historical data to predict future states of variables.
-This research area includes a broad scope of applications, such as financial investment, weather forecasting, traffic estimation, and health management Bi et al.
-(2023); Gao et al.
+(a) Performance comparison
+(b) Training Loss
+Figure 1: Two experiental findings on the ETTh1 dataset.
+Subfigure (a) shows the univariate forecasting results of KANs with different base functions, where results are minmax-normalized and 0 means the minimum MSE in the row.
+Subfigure (b) reports the training loss of one/three layers of Linear/KAN models.
+The multivariate time series forecasting (MTSF) task aims to predict future states of variables from input historical data.
+Related technologies have been applied to a broad scope of fields, such as financial investment, weather forecasting, and traffic management Gao et al.
 (2023); Savcisens et al.
 (2024); Han et al.
-(2024a).
-The machine learning community’s progress has long inspired time series forecasting technology: the popularity of early statistical learning methods gave rise to SVR and ARMIA, while the development of deep learning introduced MLP and Transformer into time series forecasting.
-At present, various time series forecasting methods cover almost all deep learning network architectures, such as RNN, CNN, Transformer, and MLP Nie et al.
-(2023); Wu et al.
-(2023); Han et al.
-(2024b).
+(2024).
+Time series forecasting methods have long been inspired by advances in the machine learning community.
+The popularity of early statistical learning methods gave rise to SVR and ARMIA Holt (2004); recurrent neural networks (RNN), which are specially designed to process sequence data, are naturally applied to time series analysis Hewamalage et al.
+(2021); TFT Lim et al.
+(2021) and Informer Zhou et al.
+(2021) promote the application of transformer architecture in time series prediction; LSTF-Linear Zeng et al.
+(2023) rethinks the value of Linear.
+At present, various deep learning network architectures, such as RNN, CNN, Transformer, and MLP, are widely used in time series forecasting Wang et al.
+(2023); Zhou et al.
+(2024); Jia et al.
+(2024).
 The forecasting models derived from different network architectures have their own advantages in forecasting performance, running speed, and resource usage.
-Although deep learning-based models have made notable progress in time series forecasting, there are still several challenges.
-The universal approximation theorem (UAT), which is the mathematical foundation of most mainstream forecasting models, cannot provide a guarantee on the necessary network sizes (depths and widths) to approximate a predetermined continuous function with specific accuracy.
-And this theory can only achieve an approximation rather than a representation.
-The limitations of UAT have become the sword of Damocles hanging over time series forecasting.
-Furthermore, the prediction mechanism of existing models is black-box, resulting in a lack of interpretability.
-These nontransparent methods are suspected of being suitable for tasks that require a low tolerance for errors, such as medicine, law and finance.
-Kolmogorov-Arnold Network (KAN) Liu et al.
-(2024a), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a novel approach to solving the above challenges.
-On the one hand, KART proves that a multivariate continuous function can be represented as a combination of finite univariate continuous functions.
-This theorem establishes the relationship between network size and input shape under the premise of representation.
-On the other hand, KAN offers a pruning strategy that simplifies the trained KAN into a set of symbolic functions, enabling the analysis of specific modules’ mechanisms, thereby significantly enhancing the network’s interpretability.
-In addition, KAN’s function fitting idea is consistent with the properties of time series, such as periodicity and trend, which is conducive to embedding prior knowledge into the network structure and improving the performance of the network.
-Despite being a relatively recent proposal, KAN, which employs a trainable 1D B-spline functions to convert incoming signals, has already sparked numerous efforts to improve or broaden its capabilities.
-Some studies propose KAN’s variants which replace the B-splines with Chebyshev polynomials SS (2024), wavelet functions Bozorgasl and Chen (2024), Jacobi polynomials Aghaei (2024), ReLU functions Qiu et al.
-(2024), etc., to accelerate training speed and improve network performance.
-Other studies introduce KAN with existing popular network structures for various applications.
-For example, ConvKAN Bodner et al.
-(2024) and GraphKAN Zhang and Zhang (2024); Xu et al.
-(2024) are proposed for image processing and graph processing.
-In summary, KANs have been extensively empirically studied in vision and language Azam and Akhtar (2024); Yu et al.
-(2024).
-However, existing studies lack a KAN-based model that considers time series domain knowledge, making it impossible to verify whether KAN is effective in time series forecasting.
-To this end, we aim to propose a KAN-based model for the time series forecasting task and evaluate its effectiveness from four perspectives: performance, integration, speed, and interpretability.
-First, we propose the Reversible Mixture of KAN Experts model (RMoK), a KAN-based time series forecasting model that uses multiple KAN variants as experts and a gating network to adaptively assign variables to specific experts for prediction.
-RMoK is implemented as a single-layer network because we hope that it can have similar performance and better interpretability than existing methods.
-Then, we use a unified training and evaluation setting to compare the performance of RMoK and current popular baselines on seven real-world datasets.
-The experimental results show that RMoK achieves state-of-the-art (SOTA) performance in most cases.
-Subsequently, we conduct a comprehensive empirical study on KAN-based models, including the comparison between KAN and Linear, the effect of integrating KANs with the Transformer, and the speed of the KAN-based models.
-Finally, we discuss the interpretability of RMoK using the example of weather prediction.
-We visualize the weights of temporal features at different time steps in KAN and find the correlation between the weight distribution and the periodicity of the data.
-To sum up, the contributions of this work include:
+However, most popular time series forecasting models are based on deep learning-based modules, whose prediction process is blackbox and prediction results lack interpretability.
+These nontransparent methods are suspected to be unsuitable for tasks with low error tolerance, such as medicine, law, and finance.
+Some traditional, non-deep learning-based methods are interpretable, but it is difficult to achieve the expected performance.
+Fortunately, a novel kind of deep learning network, the Kolmogorov-Arnold Network (KAN) Liu et al.
+(2024b), which is based on the Kolmogorov-Arnold representation theorem (KART), has become a feasible solution to the above dilemma of choosing between performance and interpretability.
+KAN offers a pruning strategy that simplifies the trained KAN model into a set of symbolic functions, making it possible to analyze the mechanisms of the network, thereby significantly enhancing the interpretability.
+Although KAN has the advantages of both performance and interpretability in theory, some works Yu et al.
+(2024a) have pointed out that it does not work well for real-world tasks.
+A few works Genet and Inzirillo (2024); Vaca-Rubio et al.
+(2024) have tried to apply KAN to time series analysis, but they have not used the characteristics of time series data to design efficient KAN-based networks, and lack fair comparisons on common datasets.
+Existing works cannot provide a convincing conclusion on whether KAN is effective in time series forecasting.
+Therefore, we aim to provide a reliable benchmark to comprehensively evaluate the effectiveness of the KAN in multivariate time series forecasting tasks for the first time.
+When applying KAN to time series forecasting, we encountered two challenges, as following.
+One is how to select the most suitable KAN.
+The KAN proposed by Ziming Liu et al.
+is not only a model but also a general function-fitting framework.
+Subsequent studies SS (2024); Bozorgasl and Chen (2024); Aghaei (2024) use different base functions to replace the B-spline function for better performance on specific tasks.
+We report the experimental results about five KAN’s variants (KANs), each with a special base function, on seven univariates of the ETTh1 dataset, as illustrated in Figure 1(a).
+This makes it cumbersome to determine the appropriate KAN for achieving the best performance.
+The other is how to train the deeper KAN.
+We build a single-layer KAN model that can achieve similar performance with a single-layer Linear model.
+But as the layer of the KAN model increases, the performance drops significantly.
+We visualize the values of loss in 10 times repeated trainings in Figure 1(b) and find that the initial loss of KAN is larger than that of Linear, and the loss of the 3-layer KAN model becomes significantly larger and more unstable.
+To address the above challenges, we first apply the mixture-of-experts technique to KAN and obtain the Mixture-of-KAN (MoK) layer, which uses a gating network to adaptively assign the input to the matching KAN layers with special base functions according to the variable dimensions.
+This structure aggregates the advantages of KAN’s variants in different cases.
+Then, we build a simple yet efficient Multi-layer Mixture-of-KAN network (MMK) based on MoK.
+By using modules that do not compromise the ability to be converted into symbolic functions, MMK achieves excellent performance while retaining the interpretability from KAN.
+Next, we analyze the phenomenon of abnormal training loss and use some training strategies to improve the performance of the proposed MMK.
+Finally, we evaluate the effectiveness of KANs in multivariate time series forecasting from four perspectives: performance, integration, speed, and interpretability.
+We compare MMK with many popular time series forecasting models on seven commonly used datasets, and experimental results show that MMK achieves state-of-the-art performance in most cases.
+We integrate MoK into recently proposed models and find that it can improve their performance.
+And we visualize the output weights of the gating network in the MoK layer and analyze the connection between KAN’s symbolic function conversion ability and domain knowledge.
+To sum up, the contributions of this paper include:
   • To the best of our knowledge, this is the first work that comprehensively discusses the effectiveness of the booming KANs for time series forecasting.
-  • To validate our claims, we propose the Reversible Mixture of KAN Experts model, which uses a single layer of the mixture of KAN experts to keep a balance between performance and interpretability.
-  • We fairly compare the performance between RMoK and baselines on seven real-world datasets, and the experimental results show that RMoK achieves the best performance in most cases.
-And we also conduct a comprehensive empirical study on KAN-based models about integration and speed.
-  • We mine the relationship between time feature weights and data periodicity through visualization, which roughly explains the mechanism of RMoK.
-In summary, compared with the baselines in terms of performance, integration, speed, and interpretability, we conclude that KAN is effective in time series forecasting.
-2 Problem Defintion
-In multivariate time series forecasting, given historical data $\mathcal{X}=[X_{1},\cdots,X_{T}]\in\mathbb{R}^{T\times C}$, where $T$ is the time steps of historical data and $C$ is the number of variates.
-The time series forecasting task is to predict $\mathcal{Y}=[X_{T+1},\cdots,X_{T+P}]\in\mathbb{R}^{P\times C}$ during future $P$ time steps.
+  • To validate our claims, we propose an interpretable Multi-layer Mixture-of-KAN network (MMK), which uses the mixture-of-KAN layer to aggregate the advantages of KAN’s variants.
+And we use some training strategies to improve performance.
+  • We fairly compare MMK with many baselines on seven real-world datasets, and the experimental results show that MMK achieves the competitive performance.
+In summary, we can conclude that KANs are effective in multivariate time series forecasting.
 2 Related Work
-3.1 Time-Series Forecasting Models
-Figure 1: The computational process of Linear and KAN layers under a certain input and output dimension.
-Although Transformer-based methods have almost become the standard in CV and NLP, various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently.
+2.1 Multivariate Time-Series Forecasting
+Various network architectures (such as Transformer, CNN, and MLP) are competing in time series forecasting recently.
 Transformer-based time series forecasting models have strong performance but high time and memory complexity.
@@
 Pyraformer Liu et al.
-(2022a) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$.
+(2022b) utilizes the pyramid attention mechanism to capture hierarchical multi-scale time dependencies with a time and memory complexity of $\mathcal{O}(T)$.
 PatchTST Nie et al.
@@
 Convolutional neural networks (CNNs) are frequently used in time series forecasting models in the form of 1D convolution, such as ModernTCN donghao and wang xue (2024) and SCINet Liu et al.
-(2022b), but TimesNet Wu et al.
+(2022a), but TimesNet Wu et al.
 (2023) takes a different approach by converting the 1D time series into a 2D matrix through the Fourier transform and then using 2D convolution for prediction.
-In addition to the above four types of network structures, there is also time series forecasting work based on new network architectures such as Mamba Dao and Gu (2024); Wang et al.
-(2024) and RWKV Peng et al.
-(2023); Hou and Yu (2024).
 2.2 Kolmogorov-Arnold Network
 The Kolmogorov-Arnold representation theorem (KART) is the mathematical foundation of the Kolmogorov-Arnold Network (KAN) Liu et al.
-(2024a), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem.
-We show the different between KAN and MLP in Figure 1.
-Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers KAN network can be represented as:
+(2024b), which makes KAN more fitting and interpretable than Multi-Layer Perceptrons (MLP) based on the universal approximation theorem.
+Given an input tensor $\mathbf{x}\in\mathbb{R}^{n_{0}}$, the structure of $L$ layers in a KAN network can be represented as:
 $\operatorname{KAN}(\mathbf{x})=\left(\boldsymbol{\Phi}_{L}\circ\boldsymbol{\Phi}_{L-1}\circ\cdots\circ\boldsymbol{\Phi}_{2}\circ\boldsymbol{\Phi}_{1}\right)\mathbf{x},$ (1)
@@
 $\mathbf{x}_{l,j}=\sum_{i=1}^{n_{l-1}}\phi_{l-1,j,i}\left(\mathbf{x}_{l-1,i}\right),\quad j=1,\cdots,n_{l},$ (2)
-where $\phi$ consists two parts: spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
-$\phi(x)=w_{b}\mathrm{SiLU}(x)+w_{s}\mathrm{Spline}(x),$ (3)
+where $\phi$ consists two parts: the spline function and the residual activation function with learnable parameters $w_{b},w_{s}$:
+$\phi(x)=w_{a}\mathrm{SiLU}(x)+w_{b}\mathrm{Spline}(x),$ (3)
 where $\mathrm{Spline}(\cdot)$ is a linear combination of B-spline functions $\mathrm{Spline}(x)=\sum_{i}c_{i}B_{i}(x)$.
-Recent studies have proposed a large number of KAN variants to expand its application scope and improve its performance.
-Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials SS (2024) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields.
-Some work extends KAN to the fields of vision and graphics and proposes ConvKAN and GraphKAN Zhang and Zhang (2024); Bodner et al.
-(2024).
-Few works try to apply KAN to time series, but there is a lack of sufficient theoretical and experimental analysis Vaca-Rubio et al.
-(2024).
-This paper aims to analyze the effectiveness of KAN and its variants in time series forecasting, we use XKAN to describe these methods uniformly (for example, TaylorKAN is KAN with Taylor polynomials).
-4 RMoK
-4.1 Mixture of KAN Experts Layer
-Figure 2: The structure of RMoK and MoK layer.
-Time series in real-world scenarios frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time.
+Recent studies have proposed many variants of KAN to expand its application scope and improve its performance.
+Some work uses functions such as wavelet functions Bozorgasl and Chen (2024), Taylor polynomials Yu et al.
+(2024b) and Jacobi polynomials Aghaei (2024) instead of spline functions to improve the performance of KAN in different fields.
+We can unify these KANs using the generalized version of Equation 3 as:
+$\phi(x)=w_{a}\mathbb{A}(x)+w_{b}\mathbb{F}(x).$ (4)
+3 Multi-layer Mixture-of-KAN Network
+3.1 Problem Defintion
+In multivariate time series forecasting, given historical data $\mathcal{X}=[X_{1},\cdots,X_{T}]\in\mathbb{R}^{T\times C}$, where $T$ is the time steps of historical data and $C$ is the number of variates.
+The time series forecasting task is to predict $\mathcal{Y}=[X_{T+1},\cdots,X_{T+P}]\in\mathbb{R}^{P\times C}$ during future $P$ time steps.
+3.2 Mixture-of-KAN Layer
+Real-world time series data frequently exhibit non-stationarity, with their statistical properties (such as mean and variance) varying over time.
 Moreover, there are significant distribution discrepancies between variables in multivariate time series.
 This poses a significant challenge to time series forecasting techniques, inevitably impacting the KAN-based methods as well.
-Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing Linear and Transformer-based networks: KAN has many variants using different spline functions.
-Considering that the special spline function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
-Following this idea, we propose the mixture of KAN experts (MoK) layer, which is a combination of KAN and mixture of experts (MoE).
-MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data.
-KAN and its variants only differ in the spline function in Equation 3, so we use $\mathcal{K}(\cdot)$ to represent these methods uniformly in this paper.
-Our proposed MoK layer with $N$ experts can be simple formed as:
+Fortunately, when it comes to dealing with distribution shift, KAN has a unique characteristic compared to existing methods: it has many variants using different base functions.
+Considering that the special base function may be suitable for modeling certain data distributions, we try to combine several KANs into a single layer and adaptively schedule them according to the input data.
+Following this idea, we propose the Mixture-of-KAN (MoK) layer, which is a combination of KAN and mixture-of-experts (MoE).
+The MoK layer uses a gating network to assign KAN layers to variables according to temporal features, where each expert is responsible for a specific part of the data.
+KAN and its variants only differ in the base function, so we use $\mathcal{K}(\cdot)$ to represent KANs uniformly in this paper.
+Our proposed MoK layer with $N$ experts can be simply formed as:
 $\mathbf{x}_{l+1}=\sum_{i=1}^{N}\mathcal{G}(\mathbf{x}_{l})_{i}\mathcal{K}_{i}(\mathbf{x}_{l}),$ (5)
 where $\mathcal{G}(\cdot)$ is a gating network.
-This mixture of experts structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
-The gating network is the key module of the MoK layer which is responsible for learning the weight of each expert from the input data.
+This mixture of expert structure can adapt to the diversity of time series, with each expert learning different parts of temporal features, thus improving performance on time series forecasting task.
+The gating network is the key module of the MoK layer, which is responsible for learning the weight of each expert from the input data.
 The softmax gating network $\mathcal{G}_{\mathrm{softmax}}$ which uses a softmax function and a learnable weight matrix $\mathbf{w}_{g}$ to schedule input data as:
@@
 (2017) which only activates the best matching top-$k$ experts.
-It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retains experts with the highest $k$ values.
+It adds Gaussian noise to input time series by $\mathbf{w}_{\mathrm{noise}}$, and uses $\mathrm{KeepTopK}$ operation to retain experts with the highest $k$ values.
 The sparse gating network can be described as:
@@
 where $\mathrm{Norm}(\cdot)$ is standard normalization.
-Table 1: The statistics of the seven TSF datasets.
+Figure 2: The architecture of MMK. Subfigure (a) is the structure of the mixture-of-KAN (MoK) layer; subfigure (b) is the pre-sample initialization strategy for the last KAN layer.
+3.3 Network Architecture
+The interpretability of KAN comes from the fact that the trained KAN model can be converted into a combination of many simple symbolic functions.
+In order to maintain this advantage, we use layers or operators with the same properties to build our KAN-based forecasting model.
+RevIN
+RevIN Kim et al.
+(2022) includes symmetrically structured normalization $\text{RevIN}^{+}$ and denormalization operations $\text{RevIN}^{-}$ in the variable dimension.
+Before sending data $X$ to the backbone network, $\text{RevIN}^{+}$ uses a learnable affine transformation to normalize the input time series of each variable.
+After the last MoK layer, outputs are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ operation.
+Each learnable parameter of RevIN corresponds to the statistical value of the time series variable, and it can be viewed as a symbolic function in form, so it does not harm the interpretability of the overall network.
+MoK Block
+Inspired by the ResNet He et al.
+(2016), we design a simple MoKBlock, which consists of a MoK layer, BatchNorm, and Dropout in sequence, and adds a residual structure.
+By stacking several MoKBlocks, MMK can obtain the stronger feature extraction capability.
+$\operatorname{MoKBlock}(\boldsymbol{h})=\operatorname{Dropout}(\operatorname{BN}(\boldsymbol{h}+\operatorname{MoK}(\boldsymbol{h}))).$ (9)
+Loss Function
+During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts.
+Following the previous work Shazeer et al.
+(2017), we add a load balancing loss function to encourage experts have equal importance.
+First, we count the weight of experts as loads, and calculate the square of the coefficient of variation ($\operatorname{CV}(\cdot)$) of the load values as additional loss:
+$\mathcal{L}_{\text{load-balancing}}=\operatorname{CV}(\text{loads})^{2}.$ (10)
+With this loss function, all experts are encouraged to have equal importance in the training stage. The total loss function is the sum of the prediction loss and load balancing loss:
+$\mathcal{L}=\mathcal{L}_{\text{pred}}+\mathcal{L}_{\text{load-balancing}},$ (11)
+where $\text{where }\mathcal{L}_{\text{pred}}=\operatorname{MSE}(\mathcal{Y},\hat{\mathcal{Y}}).$
+3.4 Training Strategy
+Although our proposed MMK is carefully designed, the performance of multiple-layer MMK is not as good as that of a single-layer MMK.
+From the visualization of the training loss, we infer that the main issue in training deep MMK is that the initial loss is too large and unstable.
+In the original KAN paper, $w_{a}$ is initialized according to the Xavier initialization and $w_{b}$ is initialized to 1 to guarantee variance-preserving Glorot and Bengio (2010); He et al.
+(2015); Yang and Wang (2024) with the assumption that all $x_{i}$ in $x$ are mutually independent.
+However, this initialization method is unstable when the number of network layers increases, because its premise, the mathematical expectation remains 0, is heavily dependent on the distribution of input data.
+We compare expectations of KAN and Linear in Equation 12:
+$\mathbb{E}(\boldsymbol{y})^{\text{KAN}}$ $=\mathbb{E}[\phi(\boldsymbol{x})]$ (12)
+$=\mathbb{E}[\boldsymbol{w}_{a}]\mathbb{E}[\mathbb{A}(\boldsymbol{x})]+\mathbb{E}[\boldsymbol{w}_{b}]\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
+$=\mathbb{E}[\mathbb{F}(\boldsymbol{x})],$
+$\mathbb{E}(\boldsymbol{y})^{\text{Linear}}$ $=\mathbb{E}[\boldsymbol{w}\boldsymbol{x}]=\mathbb{E}[\boldsymbol{w}]\mathbb{E}[\boldsymbol{x}]=0.$
+The Linear layer keeps $\mathbb{E}=0$ by initializing $\boldsymbol{w}=0$, while KAN keeps $\mathbb{E}=0$ by supposing $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$.
+Since the real-world time series data has a biased distribution and non-negligible number of outliers, it is difficult to maintain $\mathbb{E}[\mathbb{F}(\boldsymbol{x})]=0$.
+Therefore, we propose a pre-sampling initialization strategy: before training starts, we forward all training samples to the model, and record the $\operatorname{Var}[\boldsymbol{x}]$ and $\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]$ to initialize the $\boldsymbol{w}_{b}$ with $\mathcal{N}(0,\frac{\operatorname{Var}[\mathbb{F}(\boldsymbol{x})]}{\operatorname{Var}[\boldsymbol{x}]})$.
 Dataset Variates Timesteps Granularity
-ETTh1,h2 7 17,420 1 hour
-ETTm1,m2 7 69,680 15 min
+ETTh1/h2 7 17,420 1 hour
+ETTm1/m2 7 69,680 15 min
 ECL 321 26,304 1 hour
@@
 Weather 21 52,696 10 min
-Table 2: Overall performance. Both the best results of KAN-based models and baselines are highlighted in bold, and the best results of all methods are marked in red.
-Dataset P KAN-based Transformer-based CNN-based Linear-based
-RMoK-S RMoK-B PatchTST FEDFormer TimesNet SCINet RLinear DLinear
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.326 0.360 0.320 0.358 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.355 0.376 0.345 0.372
-192 0.367 0.382 0.364 0.383 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.391 0.392 0.380 0.389
-336 0.400 0.404 0.395 0.405 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.424 0.415 0.413 0.413
-720 0.462 0.439 0.457 0.440 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.450 0.474 0.453
-avg 0.389 0.396 0.384 0.397 0.387 0.400 0.448 0.452 0.400 0.406 0.485 0.481 0.414 0.407 0.403 0.407
-ETTm2 96 0.176 0.261 0.176 0.261 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.182 0.265 0.193 0.292
-192 0.244 0.305 0.240 0.302 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.246 0.304 0.284 0.362
-336 0.306 0.346 0.299 0.342 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.307 0.342 0.369 0.427
-720 0.405 0.404 0.397 0.401 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.407 0.398 0.554 0.522
-avg 0.283 0.329 0.278 0.326 0.281 0.326 0.305 0.349 0.291 0.333 0.571 0.537 0.286 0.327 0.350 0.401
-ETTh1 96 0.382 0.396 0.374 0.397 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.386 0.395 0.386 0.400
-192 0.430 0.426 0.419 0.429 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.437 0.424 0.437 0.432
-336 0.468 0.443 0.461 0.450 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.479 0.446 0.481 0.459
-720 0.450 0.458 0.474 0.467 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.481 0.470 0.519 0.516
-avg 0.433 0.431 0.432 0.436 0.469 0.454 0.440 0.460 0.458 0.450 0.747 0.647 0.446 0.434 0.456 0.452
-ETTh2 96 0.313 0.357 0.301 0.353 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.288 0.338 0.333 0.387
-192 0.397 0.409 0.379 0.405 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.374 0.390 0.477 0.476
-336 0.441 0.446 0.432 0.446 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.415 0.426 0.594 0.541
-720 0.453 0.461 0.446 0.463 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.420 0.440 0.831 0.657
-avg 0.401 0.418 0.390 0.417 0.387 0.407 0.437 0.449 0.414 0.427 0.954 0.723 0.374 0.398 0.559 0.515
-ECL 96 0.187 0.273 0.178 0.267 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.201 0.281 0.197 0.282
-192 0.193 0.280 0.187 0.274 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.201 0.283 0.196 0.285
-336 0.210 0.295 0.204 0.290 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.215 0.298 0.209 0.301
-720 0.256 0.330 0.247 0.323 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.257 0.331 0.245 0.333
-avg 0.211 0.294 0.204 0.288 0.205 0.290 0.214 0.327 0.192 0.295 0.268 0.365 0.219 0.298 0.212 0.300
-Traffic 96 0.601 0.367 0.541 0.340 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.649 0.389 0.650 0.396
-192 0.576 0.354 0.529 0.330 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.601 0.366 0.598 0.370
-336 0.588 0.359 0.545 0.334 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.609 0.369 0.605 0.373
-720 0.625 0.376 0.580 0.351 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.647 0.387 0.645 0.394
-avg 0.597 0.364 0.549 0.339 0.481 0.304 0.610 0.376 0.620 0.336 0.804 0.509 0.626 0.378 0.625 0.383
-Weather 96 0.175 0.225 0.171 0.221 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.192 0.232 0.196 0.255
-192 0.224 0.267 0.220 0.263 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.240 0.271 0.237 0.296
-336 0.284 0.308 0.277 0.302 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.292 0.307 0.283 0.335
-720 0.366 0.358 0.360 0.354 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.364 0.353 0.345 0.381
-avg 0.262 0.290 0.257 0.285 0.259 0.281 0.309 0.360 0.259 0.287 0.292 0.363 0.272 0.291 0.265 0.317
-Top-1 Count 31 23 7 12
-Table 3: The performance of KAN-based models vs. Linear-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
+Table 1: The statistics of the used datasets.
+MMK iTransformer PatchTST FEDformer TimesNet SCINet TiDE DLinear
+Models MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+96 0.374 0.397 0.386 0.405 0.414 0.419 0.376 0.419 0.384 0.402 0.654 0.599 0.479 0.464 0.386 0.400
+192 0.419 0.429 0.441 0.436 0.460 0.445 0.420 0.448 0.436 0.429 0.719 0.631 0.525 0.492 0.437 0.432
+336 0.461 0.450 0.487 0.458 0.501 0.466 0.459 0.465 0.491 0.469 0.778 0.659 0.565 0.515 0.481 0.459
+ETTh1 720 0.474 0.467 0.503 0.491 0.500 0.488 0.506 0.507 0.521 0.500 0.836 0.699 0.594 0.558 0.519 0.516
+96 0.301 0.353 0.297 0.349 0.302 0.348 0.358 0.397 0.340 0.374 0.707 0.621 0.400 0.440 0.333 0.387
+192 0.379 0.405 0.380 0.400 0.388 0.400 0.429 0.439 0.402 0.414 0.860 0.689 0.528 0.509 0.477 0.476
+336 0.432 0.446 0.428 0.432 0.426 0.433 0.496 0.487 0.452 0.452 1.000 0.744 0.643 0.571 0.594 0.541
+ETTh2 720 0.446 0.463 0.427 0.445 0.431 0.446 0.463 0.474 0.462 0.468 1.249 0.838 0.874 0.679 0.831 0.657
+96 0.320 0.358 0.334 0.368 0.329 0.367 0.379 0.419 0.338 0.375 0.418 0.438 0.364 0.387 0.345 0.372
+192 0.364 0.383 0.377 0.391 0.367 0.385 0.426 0.441 0.374 0.387 0.439 0.450 0.398 0.404 0.380 0.389
+336 0.395 0.405 0.426 0.420 0.399 0.410 0.426 0.441 0.410 0.411 0.490 0.485 0.428 0.425 0.413 0.413
+ETTm1 720 0.457 0.440 0.491 0.459 0.454 0.439 0.543 0.490 0.478 0.450 0.595 0.550 0.487 0.461 0.474 0.453
+96 0.176 0.261 0.180 0.264 0.175 0.259 0.203 0.287 0.187 0.267 0.286 0.377 0.207 0.305 0.193 0.292
+192 0.240 0.302 0.250 0.309 0.241 0.302 0.269 0.328 0.249 0.309 0.399 0.445 0.290 0.364 0.284 0.362
+336 0.299 0.342 0.311 0.348 0.305 0.343 0.325 0.366 0.321 0.351 0.637 0.591 0.377 0.422 0.369 0.427
+ETTm2 720 0.397 0.401 0.412 0.407 0.402 0.400 0.421 0.415 0.408 0.403 0.960 0.735 0.558 0.524 0.554 0.522
+96 0.166 0.256 0.148 0.240 0.181 0.270 0.193 0.308 0.168 0.272 0.247 0.345 0.237 0.329 0.197 0.282
+192 0.187 0.274 0.162 0.253 0.188 0.274 0.201 0.315 0.184 0.289 0.257 0.355 0.236 0.330 0.196 0.285
+336 0.204 0.290 0.178 0.269 0.204 0.293 0.214 0.329 0.198 0.300 0.269 0.369 0.249 0.344 0.209 0.301
+ECL 720 0.247 0.323 0.225 0.317 0.246 0.324 0.246 0.355 0.220 0.320 0.299 0.390 0.284 0.373 0.245 0.333
+96 0.511 0.324 0.395 0.268 0.462 0.295 0.587 0.366 0.593 0.321 0.788 0.499 0.805 0.493 0.650 0.396
+192 0.529 0.330 0.417 0.276 0.466 0.296 0.604 0.373 0.617 0.336 0.789 0.505 0.756 0.474 0.598 0.370
+336 0.545 0.334 0.433 0.283 0.482 0.304 0.621 0.383 0.629 0.336 0.797 0.508 0.762 0.477 0.605 0.373
+Traffic 720 0.580 0.351 0.467 0.302 0.514 0.322 0.626 0.382 0.640 0.350 0.841 0.523 0.719 0.449 0.645 0.394
+96 0.164 0.210 0.174 0.214 0.177 0.218 0.217 0.296 0.172 0.220 0.221 0.306 0.202 0.261 0.196 0.255
+192 0.210 0.251 0.221 0.254 0.225 0.259 0.276 0.336 0.219 0.261 0.261 0.340 0.242 0.298 0.237 0.296
+336 0.265 0.290 0.278 0.296 0.278 0.297 0.339 0.380 0.280 0.306 0.309 0.378 0.287 0.335 0.283 0.335
+Weather 720 0.343 0.342 0.358 0.349 0.354 0.348 0.403 0.428 0.365 0.359 0.377 0.427 0.351 0.386 0.345 0.381
+Rank 1 27 21 8 1 2 0 0 0
+Table 2: Multivariate time series forecasting results.
+Forecasting horizons $P$ is 96, 192, 336, 720, and input length $T$ is set as 96.
+A lower value indicates better performance.
+The best results are marked in red.
+4 Experiments
+In this section, we experimentally validate the effectiveness of KAN on seven public benchmarks.
+Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, integrating the KANs into other advanced models, and the interpretability of MMK.
 with MoE w/o MoE
-RMoK-S RMoL-S RWavKAN RTaylorKAN RLinear
+MoK MoL WavKAN TaylorKAN Linear
 Dataset P MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 Weather avg 0.262 0.290 0.263 0.289 0.264 0.291 0.269 0.294 0.272 0.291
-Table 4: The performance of KAN within iTransformer.
-The best results of all methods are highlighted in bold.
-KAN, JacobKAN, and TaylorKAN are KANs with B-splines functions, Jacobi polynomials, and Taylor polynomials.
-Dataset P iTransformer +KAN +JacobKAN +TaylorKAN +MoK (Ours)
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.341 0.373 0.689 0.547 0.340 0.373 0.331 0.365 0.329 0.364
-192 0.388 0.397 0.707 0.556 0.382 0.395 0.379 0.392 0.376 0.390
-336 0.426 0.420 0.720 0.565 0.418 0.419 0.417 0.416 0.412 0.415
-720 0.498 0.458 0.746 0.582 0.487 0.456 0.484 0.450 0.478 0.453
-avg 0.413 0.412 0.716 0.562 0.407 0.411 0.403 0.406 0.399 0.406
-ETTh1 96 0.390 0.404 0.696 0.557 0.401 0.414 0.397 0.409 0.393 0.409
-192 0.445 0.436 0.716 0.572 0.454 0.445 0.443 0.436 0.442 0.441
-336 0.489 0.457 0.721 0.583 0.499 0.469 0.491 0.461 0.476 0.458
-720 0.505 0.487 0.731 0.607 0.521 0.499 0.510 0.488 0.494 0.485
-avg 0.457 0.446 0.716 0.580 0.469 0.457 0.460 0.448 0.451 0.448
-4.2 Reversible Mixture of KAN Experts Model
-We can get a sophisticated KAN-based model by stacking multiple KANs or replacing the linear layers of the existing models with KANs.
-However, we try to design a simple KAN-based model that is easy to analyze while achieving comparable performance to the most state-of-the-art time series forecasting methods.
-Inspired by several successful single-layer methods Li et al.
-(2023); Zeng et al.
-(2023), we propose a simple, effective and interpretable KAN-based model, Reversible Mixture of KAN Experts Network (RMoK), which uses RevIN Kim et al.
-(2022) and single MoK layer, as shown in Figure 2.
-First, $\text{RevIN}^{+}$ (the normalization operation of RevIN) uses a learnable affine transformation to normalize the input time series of each variable.
-Then, the MoK layer obtains the prediction results based on the normalized time series features.
-Finally, the prediction results are denormalized to the original distribution space using the same affine transformation parameters in the first step by $\text{RevIN}^{-}$ (the denormalization operation of RevIN).
-During the training stage, the gating network has a tendency to reach a winner-take-all state where it always gives large weights for the same few experts.
-Following the previous work Shazeer et al.
-(2017), we add an load balancing loss function to encourage experts have equal importance.
-First, we count the weight of experts as loads, and calcuate the square of the coefficient of variation of the load values as additional loss:
-$\mathrm{L}_{\mathrm{load-balancing}}=\mathrm{CV}(\mathrm{loads})^{2}.$ (8)
-And the total loss function is the sum of the prediction loss and load balancing loss with weight $w_{l}$:
-$\mathrm{L}=\mathrm{MSE}(\mathcal{Y},\hat{\mathcal{Y}})+w_{l}\cdot\mathrm{L}_{\mathrm{load-balancing}}.$ (9)
-5 Experiments
-In this section, we experimentally validate the effectiveness of our proposed KAN-based models, RMoK, on various time series forecasting benchmarks.
-Specifically, we conduct extensive experiments, including performance comparison, running speed comparison, the impact of integrating the KAN and MoK layers into other network structures (Transformer), and the interpretability of RMoK.
+Table 3: The performance of KAN-based models vs. Linear-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
 4.1 Experimental Settings
-Dataset.
-We conduct extensive experiments on seven widely-used real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al.
+Datasets
+We conduct extensive experiments on seven widely-used datasets, including ETT(h1, h2, m1, m2) Zhou et al.
 (2021), ECL, Traffic and Weather Lai et al.
@@
 (2023), where the training, validation, and testing sets are divided according to chronological order.
-Evaluation Metrics
-Following previous works Zhou et al. (2021); Wu et al. (2023), we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
+Settings
+We implement our method and baselines in a unify code library with PyTorch  [^1] https://github.com/pytorch/pytorch and Pytorch-lightning  [^2] https://github.com/Lightning-AI/pytorch-lightning , and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs.
+In order to prevent the ‘Drop Last’ Qiu et al.
+(2024) in the test stage where the samples of the last incomplete batch are discarded, we set the test batch size to 1.
+We use the Mean Squared Error (MSE) and the Mean Absolute Error (MAE) as metrics.
+We use KAN, TaylorKAN, JacobiKAN and WavKAN to construct the MoK layer.
+The code in the supplementary files provides more details.
 Baselines
-We select 6 well-acknowledged forecasting models as our baselines, including (1) Transformer-based methods: PatchTST Nie et al.
+We select seven well-known baselines, including (1) Transformer-based methods: iTransformer Liu et al.
+(2024a), PatchTST Nie et al.
 (2023) and FEDformer Zhou et al.
@@
 (2023) and SCINet Liu et al.
-(2022b); (3) Linear-based methods: RLinear Li et al.
-(2023) and DLinear Zeng et al.
+(2022a); (3) Linear-based methods: TiDE Das et al.
+and DLinear Zeng et al.
 (2023).
-Table 5: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1. Training batch size is 64 and infer batch size is 1.
-Model Param Train (it/s) Infer (it/s)
-DLinear 139 K 102.51 360.93
-RLinear 69.9 K 92.11 345.04
-iTransformer 3.6 M 51.03 188.79
-PatchTST 7.6 M 43.92 193.58
-RKAN 1.2 M 5.94 193.25
-RWKAN 277 K 70.39 308.69
-RTKAN 208 K 77.87 316.94
-RJKAN (d=4) 345 K 76.77 280.06
-RJKAN (d=6) 483 K 58.96 254.65
-RMoK-S 972 K 34.86 190.57
-5.2 Can KAN-based Models get SOTA Performance?
-In this section, we conduct extensive experiments to compare the forecasting performance of KAN-based models with advanced baselines.
-We use two versions of our proposed RMoK: RMoK-S (the small version, which has four experts), and RMoK-B (the base version, which has eight experts).
-We compare RMoKs with six popular Transformer-based, CNN-based and Linear-based baselines on seven benchmarks.
-The experimental results are shown in Table 2, where both the best results of KAN-based models and baselines are highlighted in bold and the best results of all models are marked in red.
-The results of baselines come from published papers Wu et al.
-(2023), and the results of RMoKs are average of four times experiments with fixed seed in [0, 1, 2, 3].
-Surprisingly, RMoK achieves the best results in most cases.
-Considering that RMoK is a simple single-layer method that does not model correlations among variates, this empirical finding adequately demonstrates that KAN-based models are effective in the TSF task.
-Specifically, we can divide these seven datasets into two groups based on the number of variables.
-On the four ETT subsets with fewer variables, RMoK outperforms the baselines most, and RMoK-S and RMoK-B exhibit their own strengths under different forecasting length $P$.
-On the Weather, ECL, and Traffic datasets with more variables, RMoK-B is significantly better than RMoK-S, indicating that the mixture of experts approach is suitable for dealing with a large number of variables.
-In addition, due to the intricate spatiotemporal correlation among various variables in Traffic dataset, the Transformer-based method PatchTST achieves the best results, whereas our proposed simple RMoK achieves suboptimal results that far exceeded other baselines.
-5.3 Does KAN Outperform Linear?
-In this section, we conduct ablation experiments to compare KAN-based models with Linear-based models in three time series forecasting datasets.
-For fair comparison, we replace the four KAN experts in RMoK-S with four Linear experts to obtain a Linear-based baseline with the mixture of expert structure, naming RMoL-S.
-And we replace the entire MoK layer with a single KAN or Linear to obtain RWavKAN, RTaylorKAN and RLinear to analyze the performance of KAN variants and Linear on the TSF task.
-The experimental results are reported in Table 3, where all the results are average of 4 times experiments and the best results are highlighted in bold.
+4.2 Can KANs Get the SOTA Performance?
+In this section, we compare our proposed MMK with 7 popular Transformer-based, CNN-based and Linear-based baselines on seven custom datasets to comprehensively evaluate the performance of KAN.
+The experimental results are shown in Table 2, where the best results of all models are marked in red and the second best results are marked with underlines.
+The results of baselines come from previous paper Liu et al.
+(2024a), and the results of MMK are average of four times experiments with fixed seed [0, 1, 2, 3].
+Surprisingly, MMK achieves the best performance in most cases.
+And it significantly outperforms on the Weather dataset, reducing the MSE by 5.74% on the 96-step forecasting compared to the second-best method, iTransformer.
+We infer that this improvement comes from the fact that KAN’s function-fitting ability is more suitable for processing natural science data.
+The performance on traffic data is poor because our method does not design a module to exploit the correlation between variables.
+Subsequent research can consider designing a KAN-based variable relationship mining module to better improve model performance.
+In general, a properly designed KAN-based model can achieve the state-of-the-art performance in most cases while maintaining the advantages of interpretability.
+Dataset iTransformer +KAN +MoK
+MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.390 0.404 0.397 0.409 0.393 0.409
+192 0.445 0.436 0.443 0.436 0.442 0.441
+336 0.489 0.457 0.491 0.461 0.476 0.458
+720 0.505 0.487 0.510 0.488 0.494 0.485
+avg 0.457 0.446 0.460 0.448 0.451 0.448
+ETTm1 96 0.341 0.373 0.331 0.365 0.329 0.364
+192 0.388 0.397 0.379 0.392 0.376 0.390
+336 0.426 0.420 0.417 0.416 0.412 0.415
+720 0.498 0.458 0.484 0.450 0.478 0.453
+avg 0.413 0.412 0.403 0.406 0.399 0.406
+Table 4: The performance of KAN within iTransformer. The best results of all methods are highlighted in red.
+4.3 Does KANs Outperform Linear?
+In this section, we conduct ablation experiments to compare KAN with Linear in time series forecasting tasks.
+For fair comprison, we only use one layer of forward network that means all models in this experiment only consist of RevIN and a single network layer.
+  • MoK: It uses four KAN variants (KAN, WavKAN, TaylorKAN, and JacobiKAN) as experts for forecasting.
+  • MoL: It uses four Linears as experts for forecasting.
+  • WavKAN, TaylorKAN, Linear: They use the corresponding layer for forecasting.
+The experimental results are reported in Table 3, where all the results are the average of 4 times experiments, and the best results are highlighted in red.
 We can conclude three useful empirical experiences.
@@
 Third, the performance of KAN-based models is affected by the specific function, which may be related to the intrinsic distribution of the time series.
-Figure 3: Heatmap of the expert loads in the RMoK model on the Weather dataset.
-The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all test samples.
-5.4 Can KAN be Integrated into Other Methods?
-In this section, we verify whether KAN can be integrated into the existing time series forecasting models as a plug-in to improve performance.
-We select iTransformer Liu et al.
-(2024b) as baseline and replace the linear projections of all the attention modules with different KAN variants and MoK layer.
-The experiments are conducted on the ETT datasets.
-For fair comparison, we set the same model hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy, and repeat four times to report the average results in Table 4.
-While iTransformer with various KANs performs well only in ETTm1, iTransformer with MoK achieves the best performance in most cases on both ETTh1 and ETTm1.
-These experimental findings demonstrate that MoK is a successful form for integrating KAN into Transformer-based methods.
-(a) Data visualization
-(b) The weight of each input feature.
-Figure 4: Visualization of KAN learning the periodicity in the data.
-5.5 Are KAN-based Models Efficiency?
-We report the size of model parameters, the training and infering speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720.
-We implement all methods in a unify code library with PyTorch Paszke et al.
-(2019), and the testing platform is a GPU server with NVIDIA A100 80GB GPUs.
+Through these two groups of experiments, we can conclude that KAN has competitive performance compared to Linear in time-series forecasting.
+Models ETTh1 ETTm1
+MSE MAE MSE MAE
+TimeMixer 0.375 0.400 0.320 0.357
++MoK 0.375 0.398 0.314 0.353
+Improve 0.00% 0.50% 1.88% 1.12%
+GBT 0.398 0.418 0.319 0.370
++MoK 0.396 0.417 0.311 0.370
+Improve 0.50% 0.24% 2.51% 0.00%
+Pathformer 0.382 0.400 0.316 0.346
++MoK 0.379 0.389 0.315 0.343
+Improve 0.79% 2.75% 0.32% 0.87%
+Table 5: The performance of KAN within other model.
+4.4 Can KANs be Integrated into Other Models?
+In Section 4.2, we have demonstrated that our proposed MMK has excellent performance compared to models based on other network architectures.
+In order to further explore the upper bound of KAN’s performance in time series forecasting, we integrate KAN into some existing time series forecasting models in the form of replacing the original predictor so as to make full use of the progress of advanced time series forecasting technology.
+First, we select iTransformer Liu et al.
+(2024a) as baseline and replace the last linear predictor with a KAN or MoK layer.
+For fair comparison, we set the same network hyperparameters in all experiments where hidden dim is 512 and layer number is 2, and search the best learning rate from 1e-2 to 1e-5 through a grid searching strategy.
+As shown in Table 4, iTransformer with KAN and MoK can achieve better performance in most cases.
+Then, we intergate MoK with more models Wang et al.
+(2024); Chen et al.
+(2024); Shen et al.
+(2023) in Table 5, MoK can achieve better performance in most cases.
+These experimental findings demonstrate that MoK is a powerful form for integrating KAN into existing models.
+4.5 How to Get Better MMK?
+We evaluate the impact of different modules or training strategies on the forecasting performance.
+The pre-sampling weights initialization strategy is used in all experiments to maintain the stability of training.
+As shown in Figure 3, we report the MSE of $L$=96 and $P$=96, except for ‘+Longer Input’ where $L$=192 and $P$=96.
+First, we can find that MoK and RevIN are useful to enhance the forecasting ability.
+Then, adding model depth (‘+MoK’) only will drop the performance significance.
+The MoKBlock module and the warmup strategy can improve the performance of multi-layer MMK to achieve the best results.
+We also report the result under ‘+Longer Input’ to illustrate that MMK can get enhancement from longer input.
+Figure 3: The impact of different modules or training strategies.
+4.6 Are KANs Efficiency?
+We report the size of model parameters, the training and inferring speed of KAN-based methods and baselines on ETTh1 dataset with input length 96 and prediction length 720.
 We set training batch size to 64, infer batch size to 1.
-The KAN, WKAN, TKAN, JAKN represent KANs with B-splines functions, Wavelet function (transform is mexican hat), Taylor polynomials (order=3), and Jacob polynomials (d is degree which sets to 4 or 6).
-As shown in Table 5, the running speed of KANs are affected by the specific implementation.
-KAN variant with Taylor polynomials can achieve close running efficiency to the Linear.
-With future hardware optimization, the efficiency of KANs will still be improved.
-5.6 Are KAN-based Models Interpretable?
-In this section, we aim to analyze the interpretability of RMoK for the time series forecasting task.
+As shown in Table 6, MMK has more parameters than DLinear, but significantly less than transformer-based methods like iTransformer and PatchTST.
+The running speed of MMK is slower than that of DLinear.
+We believe that as researchers continue to optimize the computational efficiency of KAN, the running speed of MMK will be improved in the future.
+Model Param Training (it/s) Infer (it/s)
+DLinear 139 K 102.51 360.93
+iTransformer 3.6 M 51.03 188.79
+PatchTST 7.6 M 43.92 193.58
+MMK 972 K 34.86 190.57
+Table 6: Comparison of model parameters and running speed of models. All results are tested with $L$=96 and $P$=720 on ETTh1.
+4.7 Case Study
+Figure 4: Heatmap of the expert loads in the MoK on the Weather dataset. The value at location xy equal to 1 means that the gating network assigns the y-th variable to the x-th expert in all samples.
+Figure 5: The weight of each input feature.
+In this section, we aim to analyze the interpretability of MMK for the time series forecasting task.
 First, we generate a heatmap to visualize the outputs of the gating network, which decide variable-expert assignment.
-This visualization demonstrates how the RMoK simplifies the multivariate time series forecasting task into multiple univariate forecasting subtasks.
-Then, we analyze what knowledge RMoK learns from real-world time-varying systems in each subtask.
-Our proposed RMoK model uses a single MoK layer, which consists of gating network and KAN experts.
 The gating network outputs the matching score between variables and experts based on the input feature, and the top-$k$ matching experts are selected for each variable to predict the future state.
-We train an RMoK-S model with 4 experts on Weather dataset with 21 variables, then count the top-1 scores of all samples in the test set and generate a heatmap in Figure 3, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all test samples.
-From this figure, although the matching score is related to the time-varying input data, the variable is still closely related to the specific expert.
-Therefore, we roughly approximate that the trained RMoK is a linear combination of several KANs, and RMoK treats the multivariate forecasting task as multiple univariate forecasting tasks.
-Although the above process is crude, this simplification helps us analyze the explainability of RMoK in complex real-world scenarios.
-After simplification, we try to analyze why KAN is effective in time series forecasting from the perspective of univariate and single-expert subtask.
-We use the temperature variable from Weather dataset which collects data every 10 minutes to train the KAN with B-splines function as forecasting model.
-We decompose and visualize data according to the trend, season, and resid terms through statsmodels  [^1] https://www.statsmodels.org/stable/index.html library.
-As shown in Figure 4, the temperature series has obvious daily periodicity, which is consistent with life experience.
-Then, we input the complete daily period (the past 144 time steps) data to predict the state of the next time step.
-We visualize the weight of each feature dimension of the trained RMoK.
-As shown in the Figure 4(b), there are three peaks in the feature weight, which are near 0, 72 and 144 time steps.
+We count the top-1 scores of all samples in the weather dataset and generate a heatmap in Figure 4, where (x, y) = 1.0 means the y-th variable matches the x-th expert in all samples.
+We can see that the variable is closely related to the specific expert, which indicates that MoK can effectively assign KANs to variables.
+Then, we use the temperature series in the weather dataset, which collects every 10 minutes to train the KAN model.
+We input one daily period (the past 144 time steps) of data to predict the state of the next time step.
+We visualize the weight of each feature dimension of the model.
+As shown in Figure 5, there are three peaks in the feature weight, which are near 0, 72, and 144 time steps.
 0 represents the temperature at the same time step of the previous day’s period, 144 represents the temperature at the adjacent moment, and 70 represents half a period.
-These three weight peaks correspond to the three zero points of the cosine function in one period.
-Finally, we conclude that RMoK can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
+We conclude that KAN can learn the periodicity of time series, which can preliminary explain its effectiveness in time series prediction.
 5 Conclusion
-This work discusses the effectiveness of KANs in time series forecasting.
-Due to various variants with different spline functions of KAN, we propose a single-layer mixture of KAN experts model (RMoK) to alleviate the distribution variation in time series.
-We experimentally compared KAN with existing baseline methods of various network architectures in terms of performance, integration, efficiency, and interpretability.
-Experimental results on seven real-world datasets show that RMoK performs in most metrics, it is sufficient to conclude that KAN and KAN-based models are effective in the time series forecasting task, and KAN can gain a place in the increasingly fierce model structure competition.
-The single-layer KAN-based model proposed in this work is only a preliminary attempt to introduce KAN into time series forecasting.
+The Kolmogorov-Arnold Network (KAN) seems to provide a new perspective for interpretable time series forecasting.
+In this paper, we evaluate the effectiveness of KANs in time-series forecasting tasks for the first time.
+First, we design the mixture-of-KAN layer, which uses a mixture-of-experts structure to assign variables to best-matched KAN experts.
+This module aggregates the advantages of various KAN’s variants.
+Then, we propose the multi-layer mixture-of-KAN network (MMK), which achieves excellent performance while retaining KAN’s ability to be transformed into a combination of symbolic functions.
+And we use some useful tricks to address the issues in the training stage.
+Finally, we compare MMK with many popular baselines on seven datasets.
+The effectiveness of KANs in multivariate time series forecasting has been proven by extensive experiments and visualizations.
+This work is only a preliminary attempt to introduce KAN into time series forecasting.
 We hope that our work can help future studies improve the performance and interpretability of KAN-based models.
-Appendix A Implementation Details
-A.1 Datasets
-We conduct extensive experiments on seven real-world datasets, including ETT(h1, h2, m1, m2) Zhou et al.
-[2021], ECL, Traffic and Weather Lai et al.
-[2018], whose detail information is shown in Table 6.
-We follow the same data processing operations used in TimesNet Wu et al.
-[2023], where the training, validation, and testing sets are divided according to chronological order.
-The datasets we used cover multiple domains (electricity, transportation, and weather), which are sufficient to verify the generalizability of our method.
-Table 6: The information of the seven time series forecasting datasets.
-Dataset Variates Timesteps Granularity Domain
-ETTh1,h2 7 17,420 1 hour Electricity
-ETTm1,m2 7 69,680 15 min Electricity
-ECL 321 26,304 1 hour Electricity
-Traffic 862 17,544 1 hour Transportation
-Weather 21 52,696 10 min Weather
-A.2 Baselines
-To comprehensively compare KAN with other baselines under different network structures, we select six well-known forecasting models as our baselines, including (1) Transformer-based methods: PatchTST and FEDformer; (2) CNN-based methods: TimesNet and SCINet; (3) Linear-based methods: RLinear and DLinear.
-We report the details of these baselines in Table 7.
-Table 7: The information of baselines.
-Name Type Cite
-PatchTST Transformer Nie et al. [2023]
-FEDformer Transformer Zhou et al. [2022]
-TimesNet CNN Wu et al. [2023]
-SCINet CNN Liu et al. [2022b]
-RLinear Linear Li et al. [2023]
-DLinear Linear Zeng et al. [2023]
-A.3 Experimental Settings
-We implement our method and baselines in a unify code library with PyTorch  [^2] https://github.com/pytorch/pytorch and Pytorch-lightning  [^3] https://github.com/Lightning-AI/pytorch-lightning (codes will be released after acceptance), and we conduct all experiments on a GPU server with NVIDIA A100 80GB GPUs.
-To prevent the sample drop in the test phase where the samples of the last batch are discarded, we set the test batch size to 1.
-Following previous works Zhou et al.
-[2021], Wu et al.
-[2023], we use Mean Squared Error (MSE) and Mean Absolute Error (MAE) to evaluate the performance of time series forecasting.
-$\operatorname{MAE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left|x_{i}-\hat{x}_{i}\right|,$ (10)
-$\operatorname{MSE}(x,\hat{x})$ $=\frac{1}{N}\sum_{i=0}^{N}\left(x_{i}-\hat{x}_{i}\right)^{2},$
-where $x_{i}$ denotes the $i$-th ground truth, $\hat{x_{i}}$ represents the $i$-th predicted values, and $N$ represents the number of testing samples.
-Appendix B Experimental Analysis
-We report hyperparameter sensitivity and performance robustness experimental results to further evaluate the effectiveness of KAN-based modes in time series forecasting.
-B.1 Hyperparameter Sensitivity
-We conduct hyperparameter experiments to compare the performance between MoK and KAN layers.
-Results are shown in Table 8, where WavKAN, TaylorKAN and JacobKAN are KANs with wavelet functions, Taylor polynomials, and Jacobi polynomials.
-The type of spline function affects KAN’s ability to model time series, but there is currently a lack of theoretical guidance on how to choose the best suitable spline function, which can only be determined through experimental methods.
-Thus, MoK becomes a best practice for KAN, which does not require a lot of experiments to select the optimal spline function.
-This cocktail-like solution achieves better performance than a single KAN layer in most cases and achieves overall performance improvements on datasets with a large number of variables (such as Weather, Traffic, ECL).
-Table 8: The performance of KAN-based models. Both the best results of models with or w/o MoE are highlighted in bold, and the best models of all methods are marked in red.
-Setting with MoE w/o MoE
-MoK-S MoK-B RWavKAN RTaylorKAN RJacobKAN
-MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTm1 96 0.326 0.360 0.320 0.358 0.341 0.370 0.340 0.368 0.353 0.376
-192 0.367 0.382 0.364 0.383 0.374 0.384 0.380 0.386 0.386 0.391
-336 0.400 0.404 0.395 0.405 0.406 0.405 0.412 0.406 0.413 0.410
-720 0.462 0.439 0.457 0.440 0.465 0.438 0.474 0.439 0.465 0.442
-avg 0.389 0.396 0.384 0.397 0.397 0.399 0.402 0.400 0.404 0.405
-ETTm2 96 0.176 0.261 0.176 0.261 0.179 0.265 0.181 0.264 0.177 0.263
-192 0.244 0.305 0.240 0.302 0.247 0.309 0.247 0.306 0.246 0.308
-336 0.306 0.346 0.299 0.342 0.311 0.349 0.308 0.345 0.309 0.348
-720 0.405 0.404 0.397 0.401 0.412 0.408 0.405 0.401 0.406 0.403
-avg 0.283 0.329 0.278 0.326 0.287 0.333 0.285 0.329 0.284 0.331
-ETTh1 96 0.382 0.396 0.374 0.397 0.396 0.402 0.388 0.398 0.389 0.411
-192 0.430 0.426 0.419 0.429 0.439 0.430 0.438 0.424 0.434 0.437
-336 0.468 0.443 0.461 0.450 0.479 0.449 0.477 0.442 0.464 0.451
-720 0.450 0.458 0.474 0.467 0.464 0.461 0.477 0.461 0.452 0.460
-avg 0.433 0.431 0.432 0.436 0.444 0.435 0.445 0.431 0.435 0.440
-ETTh2 96 0.313 0.357 0.301 0.353 0.321 0.358 0.297 0.344 0.310 0.352
-192 0.397 0.409 0.379 0.405 0.407 0.413 0.387 0.401 0.396 0.405
-336 0.441 0.446 0.432 0.446 0.452 0.450 0.434 0.441 0.437 0.441
-720 0.453 0.461 0.446 0.463 0.487 0.477 0.448 0.458 0.445 0.456
-avg 0.401 0.418 0.390 0.417 0.417 0.424 0.392 0.411 0.397 0.414
-ECL 96 0.187 0.273 0.178 0.267 0.202 0.284 0.203 0.282 0.201 0.284
-192 0.193 0.280 0.187 0.274 0.203 0.285 0.205 0.285 0.200 0.285
-336 0.210 0.295 0.204 0.290 0.218 0.299 0.221 0.300 0.216 0.301
-720 0.256 0.330 0.247 0.323 0.262 0.332 0.265 0.333 0.260 0.333
-avg 0.211 0.294 0.204 0.288 0.221 0.300 0.224 0.300 0.220 0.301
-Weather 96 0.175 0.225 0.171 0.221 0.180 0.228 0.182 0.232 0.178 0.230
-192 0.224 0.267 0.220 0.263 0.227 0.268 0.235 0.274 0.229 0.271
-336 0.284 0.308 0.277 0.302 0.284 0.308 0.291 0.312 0.288 0.310
-720 0.366 0.358 0.360 0.354 0.368 0.360 0.370 0.360 0.368 0.359
-avg 0.262 0.290 0.257 0.285 0.264 0.291 0.269 0.294 0.266 0.292